OpenStudy (anonymous):

If we decrease the time it takes for a car to travel over the same distance, this will increase the speed. have no effect on velocity or speed. decrease the speed. decrease the velocity.

OpenStudy (theeric):

Hi! We're looking at speed or velocity, distance, and time! If we're just talking about velocity, there's a relation between them that is very simple when written in math. Do you know the equation for velocity? That's a great place to start.

OpenStudy (anonymous):

decrease in speed

OpenStudy (theeric):

Why do you think that?

OpenStudy (theeric):

\(v=\dfrac{d}{t}\)

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

idk what happens next im not good at science

OpenStudy (theeric):

Well, it says the distance \(d\) is the same, and \(t\) decreases. \(v=\dfrac{d}{t}\) \(\dfrac{d}{\Tiny t}=\) Now the denominator is smaller than it was. The numerator is bigger in comparison. So, how will the velocity change? The other way to think is that you're running on a track. You go around once. Then you go around again in a shorter time. Was your velocity faster or slower on that second lap?